OpenStudy (anonymous):

Find the coordinates of the other endpoint if the midpoint is M(8, 2) and the other endpoint is P(5, 6). (6.5, 4) (21, 10) (11, -2) Help? Please

OpenStudy (karatechopper):

Alright, which formula do you think will help you solve this?

OpenStudy (anonymous):

I want to say midpoint but that's not right

OpenStudy (karatechopper):

It is right. :)

OpenStudy (karatechopper):

Okay. So you already know your midpoint. Do you realize that midpoint breaks into two separate parts in order to solve right? We solve for x, and then for y?

OpenStudy (anonymous):

Okay, so how do you break it into two parts?

OpenStudy (karatechopper):

Well let's do the x value first shall we? Midpoint formula looks like this: [(x_1 + x_2)/2 , (y_1 + y_2)/2] ----IN TOTAL But let's look at it in parts. To solve for Xm m=midpoint \[\frac{ x1+x2 }{ 2 }\] That's x-value 1 and 2 btw. So I'll show you how to set up and how to find X-value we are looking for. Let's call the point we are looking for, x-value 1. and we know that the Xm is 8. Have you followed what I am saying so far?

OpenStudy (anonymous):

Yes

OpenStudy (karatechopper):

Great! So let's plug values in that we know. \[\frac{ x1+5 }{ 2 }= 8\] --------I took the 5 from the x-value from point P and the 8 is the x-value from Midpoint.

OpenStudy (karatechopper):

Following?

OpenStudy (anonymous):

Okay, but how did you get 8?

OpenStudy (karatechopper):

Remember, Midpoint formula is used to find the midpoint. We were given the x-value of 8 as your midpoint in the question. We are using Midpoint formula in order to find your other point.

OpenStudy (karatechopper):

\[\frac{ x1+x2 }{ 2 }=Xm\]

OpenStudy (anonymous):

Oh right. Sorry I got confused

OpenStudy (karatechopper):

Refer to that if it helps you understand why I picked which points. :)

OpenStudy (karatechopper):

Alright so we can continue?

OpenStudy (anonymous):

Yes please

OpenStudy (karatechopper):

Alright, now here's the part where I want to see if you can solve for the x-value by yourself. I set up the equation for you, now please solve for x1. I will check it for you. :) \[\frac{ x1+5 }{ 2 }=8\]

OpenStudy (anonymous):

Don't you first multiply everything by 2 to get rid of the fraction?

OpenStudy (karatechopper):

Correct.

OpenStudy (anonymous):

so is it x1+10=16? is that correct so far?

OpenStudy (karatechopper):

You don't multiply the 2 by the 5. When you multiple the 2 to get rid of the fraction then it cancels from the first side leaving you with only x1+5 and the other side being 16. Got it?

OpenStudy (anonymous):

Oh. okay, yes got it

OpenStudy (anonymous):

so x1=11

OpenStudy (karatechopper):

Good! Now can you set up the equation for y-value and try to solve that? It's the same steps in plugging and solving. Try it! :)

OpenStudy (anonymous):

So the answer is 11, -2?

OpenStudy (karatechopper):

Perfect!

OpenStudy (karatechopper):

Don't forget parenthesis

OpenStudy (anonymous):

Thank you :)

OpenStudy (karatechopper):

You are welcome. :D